For a molecule A, the plot of total polarization Pm vs 1/T has slope while for molecule B, it has a finite positive slope. ChCl3 and CCl4 which one could be molecule A and why
1
Expert's answer
2020-10-16T08:24:02-0400

Solution.

"Pm = \frac{Na}{3E0}(\alpha + \frac{\mu^2}{3kT})"

For CCl4 "\mu" = 0, so

"Pm = \frac{Na \times \alpha}{3E0}"

Pm does not depend on temperature in the case when the dipole moment is 0. Therefore, the A molecule is CCl4. The B - CHCl3 Molecule.

Answer:

Therefore, the A molecule is CCl4. The B - CHCl3 Molecule.